OpenStudy (anonymous):

If the dimensions of a triangle are cut in half, then the new perimeter is one fourth of the original perimeter. True False

OpenStudy (anonymous):

Well, easy way to check this is to use a test case. Let's use a really easy triangle, an equilateral triangle with side length 2. So, the perimeter of this triangle would be:\[P_1 = 2+2+2=6\]Now, let's cut the dimensions of each side in half, so each side length is now 1. \[P_2 = 1 + 1 + 1 = 3\]So, based on that, what do you think the answer is?

OpenStudy (anonymous):

t